OpenStudy (anonymous):

Write the quadratic equation having the given numbers as solutions -3 and -1. The quadratic equation is blank=0

OpenStudy (anonymous):

(x+3)(x+1)

OpenStudy (anonymous):

x^2+4x+3=0

OpenStudy (anonymous):

Write the quadratic equation in the variable x having the given numbers as solutions. the equation is standard form, ax^2+bx+c=0

OpenStudy (anonymous):

\[x ^{2}+4x+3\]=0